OpenStudy (anonymous):

Change this radical to an algebraic expression with fractional exponents. exponet ^5 radical x exponet ^3

OpenStudy (anonymous):

\[ \sqrt[a]{x^b} =x^{a/b} \]

OpenStudy (jdoe0001):

\(\Large \bf \sqrt[5]{x^3}\quad ?\)

OpenStudy (anonymous):

\[ \sqrt[5]{x^3} = \ldots \]

OpenStudy (anonymous):

Yeah thats the problem jdoe0001

OpenStudy (jdoe0001):

anyhow as wio already pointed out \(\Large \bf \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\)

OpenStudy (anonymous):

\[ \sqrt[\color{red}a]{x^{\color{blue}b}} =x^{\color{red}a/\color{blue}b} \]\[ \sqrt[\color{red}5]{x^{\color{blue}3}} \]

OpenStudy (anonymous):

THANK YOU!!!!